OpenStudy (anonymous):

shiw that only one out of n,n+2,n+4is divisible by 3,where n is any positive integer @hartnn @Rohith15

OpenStudy (anonymous):

@jbreezy622

OpenStudy (anonymous):

If n is divisible by 3 there is nothing more to prove. Suppose that n is not divisible by 3 Then the remainder of dividing n by 3 is either 1 or 2. In the first case n+2 is divisible by 3; in the second case n+4 is divisible by 3.

OpenStudy (anonymous):

@Rohith15 cud u plzz write with the steps sir

Miracrown (miracrown):

What type of proof techniques have you been learning about in class?

OpenStudy (anonymous):

Suppose that n is divisible by 3. Then, we can write: n = 3k, for some integer k. Then, n + 2 = 3k + 2 and n + 4 = 3k + 4 are both not divisible by 3 since: (n + 2)/3 = k + 2/3 and (n + 4)/3 = k + 4/3 = (k + 1) + 1/3, which shows that n + 2 and n + 4 leave remainders of 2 and 1 respectively when divided by 3. In a similar fashion, if n + 2 is divisible by 3, then: n + 2 = 3k, for some integer k. Then, n = (n + 2) - 2 = 3k - 2 and n + 4 = (n + 2) + 2 = 3k + 2 are both not divisible by 3 (this can be shown similarly to the above). In a similar fashion, if n + 4 is divisible by 3, both n and n + 2 are not divisible by 3

OpenStudy (anonymous):

@Rohith15 didnt get it

Miracrown (miracrown):

There is a very elegant way to show this one. n + 4 is divisible by 3 if and only if n + 1 is divisible by 3

OpenStudy (anonymous):

@surd First read it fully and then tell me where you didnt understand.

Miracrown (miracrown):

So proving this problem is equivalent to showing that either n, n+1, or n+2 is divisible by 3. (n+4 and n+1 are the same with respect to divisibility by 3)

OpenStudy (anonymous):

plzz shiw statemen in some link or wirite it nsd show me plzz\

OpenStudy (anonymous):

See, if n is divisible by 3 we can write n=3k right??

Miracrown (miracrown):

 Now, showing that for three consecutive numbers is relative simple. If you divide by n, n+1, n+2, only one of them has a remainder of 0. So only is divisible by 3. That is simple conquence from that they are consecutive.

OpenStudy (anonymous):

yup

Miracrown (miracrown):

sorry, if you divide n, n+1, and n+2 BY 3, only one of them...

OpenStudy (anonymous):

I agree with you @Miracrown but you cant write it as a theorem right??

Miracrown (miracrown):



OpenStudy (anonymous):

mira do you mind showing me all the steoppd plzz

OpenStudy (anonymous):

@hartnn plzz go chek eng chat they nedd ur help

Miracrown (miracrown):

Why do think it couldn't be a theorem?

OpenStudy (anonymous):

PLZZ SHOW ALL THE STEPS BEG UUU

Miracrown (miracrown):

That's your job. I've explained you how to do it, now what's left is for you to do.

OpenStudy (anonymous):

@Miracrown plzz beg uuuuu